# Closed-Loop Neural Operator-Based Observer of Traffic Density

### v1

@@ III-C Closed-Loop Observer
Figure 2: Closed-loop observation with $\hat{\mathcal{S}}^{cl}$. The horizontal black lines indicate sensor locations where measurements are taken.
The correction operator $\mathcal{N}$ is approximated using an FNO defined on a two-dimensional domain,
@@
where $\Gamma=[\Delta t,(n_{d}+1)\Delta t]$ and the processed functions are translated in time to take values in this domain. By using an FNO processing scalar-valued functions on a two-dimensional domain (21) instead of an FNO processing vector-valued functions on a one-dimensional domain, as suggested by (20), the combined codomain dimension of the input functions is reduced from $2(n_{d}+1)$ to just two. Furthermore, $\Gamma$ reflects the choice of window size in (19). This choice enables an extension of the learning objective for open-loop prediction (7), such that $\mathcal{N}_{\psi}$ is obtained as the minimizer of
@@
where $\rho_{\mathcal{D}}[a]$ is a data-based estimate of the true density $\mathcal{G}_{1:n_{d}+1}[a]$ using samples at the sensor locations, $\{\mathcal{G}_{1:n_{d}+1}[a](\cdot)\}_{x\in\mathcal{M}}$. By comparing the objective function of $\mathcal{N}_{\psi}$ (22) with that of the open-loop predictor $\mathcal{G}_{\theta}$ (7), we observe that $\mathcal{N}_{\psi}$ is trained to refine the prediction of $\mathcal{G}_{\theta}$. This refinement is obtained by taking in measurements in a Luenberger manner through $\mathcal{E}$.
@@ IV Numerical Simulations and Discussion
Before presenting the experimental results, the implementation details are outlined. We then illustrate the observers on a single test example; analyse the performance in the noiseless, noisy, and OOD setting; and analyse the input-to-state stability.
@@ IV-A Implementation Details
The SUMO simulations are generated on a circular road of length $L=6.2$ km discretized into 123 cells $\mathcal{X}$ with six equidistant sensors $\mathcal{M}$. The simulations are run for a total time of $T=40$ minutes divided into time steps of length $\Delta t=1$ s, following [17]. The training dataset contains 20 noiseless, independent simulations per average $\rho\in[0.1,0.2,...,0.8]$ split into a total of 3360 non-overlapping input-output pairs $\big{(}\zeta_{0}^{j},[\rho^{j}_{k\Delta t}]_{k=1}^{n_{d}+1}\big{)}$. We found that $n=10$ and $n_{d}+1=100$ works well in this setting.
The test dataset consists of 10 simulations per average $\rho\in[0.3,...,0.8]$, focusing on the settings where congestion occurs. In addition, we test the observers on a noisy test set, and a noiseless OOD test set. The noisy data is produced by adding i.i.d. Gaussian noise with standard deviation $\sigma=0.1$ to the measurements. The OOD data is generated by varying parameters in SUMO to induce more traffic jams.
The approximate solution operator $\mathcal{G}_{\theta}$ and correction operator $\mathcal{N}_{\psi}$ share a similar FNO structure: the lifting function is a linear layer $P$ of output dimension 16, and the projection function $Q$ is a fully-connected neural network $Q$ with one hidden layer of size 128. In between, $\mathcal{G}_{\theta}$ has four spectral convolution layers $\mathcal{L}_{l}$ where $W_{l}$ are one-dimensional convolution layers of sizes $[24,24,32,32]$ and the Fourier expansions are truncated to $k_{\max}$ modes $[15,12,9,9]$. The correction operator $\mathcal{N}_{\psi}$ has two spectral convolution layers $\mathcal{L}_{l}$ where $W_{l}$ are two-dimensional convolution layers of sizes $[24,32]$ and the Fourier expansions are truncated to $k_{\max}$ modes $[15,9]$. For both operators, we use GELU activation functions and warp the output with sigmoid to impose $\hat{\rho}\in[0,1]$. GP regression is implemented using a prior with zero mean function and squared exponential kernel with length scale 1, and we sample once during training while taking the posterior mean function during inference.
We use the Adam optimizer and train for 500 epochs. The training and testing were run on a laptop with an Intel(R) Core(TM) i7-1370p CPU @ 1.90 GHz and 14 processing cores. The code and experiments are available on GitHub for reproduction [^1] https://github.com/aharting.
@@ IV-B Numerical Simulations and Discussion
A comparison between the open-loop observer $\hat{\mathcal{S}}^{ol}$, open-loop observer with reset $\hat{\mathcal{S}}^{ol\text{-}r}$, and closed-loop observer $\hat{\mathcal{S}}^{cl}$ applied to a single test example is shown in Fig. 3. We observe that $\hat{\mathcal{S}}^{ol}$ quickly diverges, while $\hat{\mathcal{S}}^{ol\text{-}r}$ generates disconnected, volatile predictions. In contrast, $\hat{\mathcal{S}}^{cl}$ maintains stability while producing consistent predictions over the spatiotemporal domain. This highlights the benefits of $\hat{\mathcal{S}}^{cl}$: the continuous calibration of the autoregressive inputs both improves the prediction estimates and counteracts instability. Equipped with a stable autoregressive mechanism, $\hat{\mathcal{S}}^{cl}$ enables physically coherent predictions between the sparse measurements. We continue by exploring the performance across the test set in terms of accuracy, robustness, and input-to-state stability.
Figure 3: Traffic density estimation with observer variants. The true density is shown in the top left figure, where measurement locations are indicated in black. The density estimates are shown for the open-loop observer (top right), the open-loop observer with reset (bottom left), and the closed-loop observer (bottom right).
@@
The prediction accuracy of the observers across the test set in the noiseless, noisy, and OOD settings is shown in Fig. 4. In the noiseless setting, the closed-loop observer $\hat{\mathcal{S}}^{cl}$ generally performs better than the open-loop observer $\hat{\mathcal{S}}^{ol}$ and the open-loop observer with reset $\hat{\mathcal{S}}^{ol\text{-}r}$, since the median error is lower. The poor performance of $\hat{\mathcal{S}}^{ol}$ is explained by the instability caused by disturbance in the initial state. The improved performance of $\hat{\mathcal{S}}^{cl}$ over $\hat{\mathcal{S}}^{ol\text{-}r}$ can be explained by the fact that $\hat{\mathcal{S}}^{cl}$ benefits from the full historical dataset, as previous corrections influence all subsequent predictions through the continuous chain of autoregression.
@@
Moreover, $\hat{\mathcal{S}}^{cl}$ is generally robust to noise, as the median error remains around the same level when the observer is applied to noisy data. This is not the case of $\hat{\mathcal{S}}^{ol\text{-}r}$, which sees a significant downgrade in performance. This can be explained by the high dependence on few measurements in $\hat{\mathcal{S}}^{ol\text{-}r}$, rendering it vulnerable to noise. Instead, $\hat{\mathcal{S}}^{cl}$ weighs the recent measurements against the prediction suggested by past measurements through the correction operator, resulting in a more accurate estimate. Furthermore, the accuracy of $\hat{\mathcal{S}}^{cl}$ generally remains the same in the OOD dataset, suggesting that it is robust to distribution shifts. Overall, our simulations indicate that $\hat{\mathcal{S}}^{cl}$ is both more accurate and more resilient to noise compared to $\hat{\mathcal{S}}^{ol}$ and $\hat{\mathcal{S}}^{ol\text{-}r}$, as well as robust to distributional changes.
@@ V Conclusion
In this work, we introduced a data-driven, closed-loop observer for traffic density estimation using sparse stationary sensor measurements. Our method trains a Fourier neural operator to learn macroscopic traffic flow dynamics from microscopic simulations, and integrates it with a correction operator to update predictions with real-time data. Numerical simulations show that the closed-loop observer outperforms open-loop alternatives in terms of accuracy, robustness to noise, and stability over time, and indicate that the closed-loop observer is robust to distribution shifts.
This study highlights the potential of this kind of learning-based, closed-loop observer and opens several interesting research directions. Future work includes extending the observer to mobile sensors, leveraging traffic data with varying resolution to learn the dynamics, and conducting formal analyses of convergence and stability.

### v2

@@ III-C Closed-Loop Observer
The correction operator $\mathcal{N}$ is approximated using an FNO defined on a two-dimensional domain:
@@
where $\Gamma=[\Delta t,(n_{d}+1)\Delta t]$ and the processed functions are translated in time to take values in this domain. Note that $\Gamma$ reflects the choice of window size in (28). This setting enables an extension of the learning objective for open-loop prediction (8), such that $\mathcal{N}_{\psi}$ is obtained as the minimizer of
@@
where $\rho_{\mathcal{D}}[a]$ is a data-based estimate of the true density $\mathcal{G}_{1:n_{d}+1}[a]$ using samples at the sensor locations, $\{\mathcal{G}_{1:n_{d}+1}[a](\cdot)\}_{x\in\mathcal{M}}$. By comparing the objective function of $\mathcal{N}_{\psi}$ in (31) with that of the open-loop predictor $\mathcal{G}_{\theta}$ in (8), we observe that $\mathcal{N}_{\psi}$ is trained to refine the prediction of $\mathcal{G}_{\theta}$. This refinement is obtained by taking in measurements in a Luenberger manner through $\mathcal{E}$.
@@ IV Numerical Simulations and Discussion
Numerical simulations are conducted using the open-source microscopic traffic simulator SUMO [27]. This framework integrates road network data, infrastructure like traffic lights, and stochastic demand to produce realistic vehicle flows. Following the methodological section, we focus on ring road dynamics [28]: vehicles are gradually introduced to the road during an initialization phase to reach the initial state, after which the higher-order microscopic dynamics are run in a loop. To convert the discrete distribution of vehicles to a continuous density function, the grid of vehicle-populated cells is convolved with a Gaussian filter [29].
The remainder of this section is organized as follows: we first outline implementation details, then demonstrate the observers on a single test case, evaluate their performance in noiseless, noisy, and out-of-distribution (OOD) scenarios, and finally analyze input-to-state stability.
@@ IV-A Implementation Details
The SUMO simulations are generated on a circular road of length $L=6.2$ km, discretized into 123 cells $\mathcal{X}$ with six equidistant sensors $\mathcal{M}$. The simulations are run for a total time of $T=40$ minutes divided into time steps of length $\Delta t=1$ s, following [28]. The training dataset contains 20 noiseless, independent simulations per average $\rho\in[0.1,0.2,...,0.8]$ split into a total of 3360 non-overlapping input-output pairs $\big{(}\mathrm{P}_{0}^{j},[\rho^{j}_{k\Delta t}]_{k=1}^{n_{d}+1}\big{)}$. We found that $n=10$ and $n_{d}+1=100$ works well in this setting.
The test dataset consists of 10 simulations per average $\rho\in[0.3,...,0.8]$, focusing on the settings where congestion occurs. In addition, we test the observers on a noisy test set and a noiseless OOD test set. The noisy data is produced by adding i.i.d. Gaussian noise with standard deviation $\sigma=0.1$ to the measurements. The OOD data is generated by varying parameters in SUMO to induce more traffic jams.
The approximate solution operator $\mathcal{G}_{\theta}$ and correction operator $\mathcal{N}_{\psi}$ share a similar FNO structure: the lifting function is a linear layer $P$ of output dimension 16, and the projection function $Q$ is a fully-connected neural network $Q$ with one hidden layer of size 128. In between, $\mathcal{G}_{\theta}$ has four spectral convolution layers $\mathcal{L}_{l}$ where $W_{l}$ are one-dimensional convolution layers of sizes $[24,24,32,32]$ and the Fourier expansions are truncated to $k_{\max}$ modes $[15,12,9,9]$. The correction operator $\mathcal{N}_{\psi}$ has two spectral convolution layers $\mathcal{L}_{l}$ where $W_{l}$ are two-dimensional convolution layers of sizes $[24,32]$ and the Fourier expansions are truncated to $k_{\max}$ modes $[15,9]$. For both operators, we use GELU activation functions and warp the output with a sigmoid to impose $\hat{\rho}\in[0,1]$. GP regression is implemented using a prior with a zero mean function and a squared exponential kernel with a length scale of 1, and we sample once during training while taking the posterior mean function during inference.
We use the Adam optimizer and train for 500 epochs. The training and testing were run on a laptop with an Intel(R) Core(TM) i7-1370p CPU @ 1.90 GHz and 14 processing cores. The code and experiments are available on GitHub for reproduction [^3] https://github.com/aharting/closed-loop-neural-operator-based-observer-of-traffic-density.
@@ IV-B Numerical Simulations and Discussion
A comparison between the open-loop observer $\hat{\mathcal{S}}^{ol}$, open-loop observer with reset $\hat{\mathcal{S}}^{ol\text{-}r}$, and closed-loop observer $\hat{\mathcal{S}}^{cl}$ applied to a single test example is shown in Fig. 7. We observe that $\hat{\mathcal{S}}^{ol}$ quickly diverges, while $\hat{\mathcal{S}}^{ol\text{-}r}$ produces disconnected predictions, with the vertical dispersion of the measurements resulting from the vertical GP interpolation between observation points. In contrast, $\hat{\mathcal{S}}^{cl}$ maintains stability while producing consistent predictions over the spatiotemporal domain. This highlights the benefits of $\hat{\mathcal{S}}^{cl}$: the continuous calibration of the autoregressive inputs both improves the prediction estimates and counteracts instability. Hence, $\hat{\mathcal{S}}^{cl}$ unites the consistent predictions of $\hat{\mathcal{S}}^{ol}$ with the online updates of $\hat{\mathcal{S}}^{ol\text{-}r}$, achieving an improved observer. We continue by exploring the performance across the test set in terms of accuracy, robustness, and input-to-state stability.
@@
The prediction accuracy of the observers across the test set in the noiseless, noisy, and OOD settings is shown in Fig. 8. In the noiseless setting, the closed-loop observer $\hat{\mathcal{S}}^{cl}$ generally performs better than the open-loop observer $\hat{\mathcal{S}}^{ol}$ and the open-loop observer with reset $\hat{\mathcal{S}}^{ol\text{-}r}$. The poor performance of $\hat{\mathcal{S}}^{ol}$ is explained by the instability caused by disturbance in the initial state. The improved performance of $\hat{\mathcal{S}}^{cl}$ over $\hat{\mathcal{S}}^{ol\text{-}r}$ can be explained by the fact that $\hat{\mathcal{S}}^{cl}$ benefits from the full historical dataset, as previous corrections influence all subsequent predictions through the continuous chain of autoregression.
Figure 7: Traffic density estimation with observer variants. The true density is shown in the top left figure, where measurement locations are indicated in black. The density estimates are shown for the open-loop observer (top right), the open-loop observer with reset (bottom left), and the closed-loop observer (bottom right). Note that the solution periodicity is a result of the ring road implementation.
@@
Moreover, $\hat{\mathcal{S}}^{cl}$ is generally robust to noise, as the median error remains around the same level when the observer is applied to noisy data. This is not the case of $\hat{\mathcal{S}}^{ol\text{-}r}$, which sees a significant downgrade in performance. This can be explained by the high dependence on a few measurements in $\hat{\mathcal{S}}^{ol\text{-}r}$, rendering it vulnerable to noise. Instead, $\hat{\mathcal{S}}^{cl}$ weighs the recent measurements against the prediction suggested by past measurements through the correction operator, resulting in a more accurate estimate. Furthermore, the accuracy of $\hat{\mathcal{S}}^{cl}$ generally remains the same in the OOD dataset, suggesting that it is robust to distribution shifts. Overall, our simulations indicate that $\hat{\mathcal{S}}^{cl}$ is both more accurate and more resilient to noise compared to $\hat{\mathcal{S}}^{ol}$ and $\hat{\mathcal{S}}^{ol\text{-}r}$, as well as robust to distributional changes.
@@ V Conclusion
In this work, we introduced a data-driven, closed-loop observer for traffic density estimation using sparse stationary sensor measurements. Our method trains a Fourier neural operator to learn macroscopic traffic flow dynamics from high-fidelity data and integrates it with a correction operator to update predictions with real-time data. Numerical simulations with SUMO show that the closed-loop observer outperforms open-loop alternatives in terms of accuracy, robustness to noise, and stability over time, and indicate that the closed-loop observer is robust to distribution shifts.
This study highlights the potential of this kind of learning-based, closed-loop observer and opens several interesting research directions. Future work includes leveraging real traffic data – possibly with varying resolution – to learn the dynamics, analyzing sensor placement for observability, and conducting formal analyses of convergence and stability.